# It Is Time To Steer: A Scalable Framework for Analysis-driven Attack Graph Generation

### v1

@@ 4.1 StatAG Validation
To validate StatAG we first present the experimental setting, and then we analyze the convergence of vulnerability features, attack path features, and stability distributions.
4.1.1 Experimental Setting
@@
We validate StatAG with an experimental setup designed for approaching the different factors that affect the AG scalability [15, 32]. It consists of synthetic networks and vulnerability inventories, in which we vary the number of hosts, vulnerabilities, network topology, and composition of vulnerabilities per host. Looking at the former, we considered configurations with up to 20 hosts and 50 vulnerabilities per host, which is the biggest configuration possible to compute all the paths (complete AG, used as ground truth) in a reasonable amount of time (in the order of hundreds of hours) [14]. We then tested $3$ network topologies, including mesh [45], random [46], and power law [47]. We considered $5$ levels of heterogeneity of vulnerabilities, varying from 0% (i.e., all vulnerabilities among the hosts are the same) to 100% (i.e., all vulnerabilities among the hosts are different) to increase the variability of vulnerability features. Thus, for each network, we evaluated 15 different configurations, which we varied in 100 statistical variations, for a total of 1500 experiments.
We use Python and the scikit-learn machine learning library [48, 49] for the implementation and we run the experiments on a Linux server with Intel(R) Xeon(R) Gold 6248 CPU 2.50GHz and 256 GB memory. All the materials, as well as the source code of all components including the code used for the experimental evaluation, are publicly available [^1] https://github.com/Ale96Pa/attack_graph_progressive.
4.1.2 Results
@@
We validate StatAG by studying the convergence of the partial AGs to the complete AG. For the sake of simplicity, in the rest of this section, we refer to the complete AG as ground truth (GT) to indicate that it corresponds to all the paths generated at the end of the classic generation process.
Figure 3: KS distance trend of partial AGs from the GT for the vulnerability features distribution.
@@
The first validation step concerns the convergence of the vulnerability feature distributions of the partial AGs to the GT. Fig. 3 reports the median KS distances during the entire execution of the approach. The convergence speed of the vulnerability feature distributions depends on the distance threshold $T$ (see Definition 6). For example, considering a good approximation as $T=0.1$ (corresponding to just 10% approximation of the GT), 9 of 10 features are statistically significant after only 100 iterations, except for the availability feature that has a KS distance of 0.2 after 100 iterations. In particular, we can notice 3 different trends in all the plots in Fig. 3. The first one coincides with early partial results, where the KS distance from the GT is too high to perform analysis with a reasonable approximation but good for initial exploration. This trend lasts about the first 100 iterations. After 100 iterations, corresponding to 9% of the complete execution, the trend is flat until 1700 iterations on average, corresponding to 77% of the complete execution. This is the interval where mature partial results appear, that provide a 10% approximation of the GT for most vulnerability features, with the worst performance reached by the availability with a 20% approximation. Finally, in the last 23% of the execution, the KS distance becomes very close to 0, corresponding to the definitive partial results. The duration of mature partial results, in contrast to early and definitive ones, can be attributed to the overall limited variability in vulnerability features according to the CVSS metrics, as documented by NIST [24]. As a consequence, the vulnerability variability is covered very quickly in just 9% of the total number of iterations.
The convergence trend of the vulnerability features to the GT shows the capability of the progressive approach to represent 90% of the vulnerability inventory after a few iterations. It is a good result for the analysis of the vulnerability inventory, but we need further investigation that concerns the attack path analysis. To this aim, the next validation step concerns the convergence trend of the attack path features of the partial AGs to the GT ones. Fig. 4 reports the distribution trend of the KS distances between the partial AGs and the GT for the attack path features, i.e., likelihood and impact.
Figure 4: KS distance trend of partial AGs from the GT for the attack path features distribution.
@@
The trends in Fig. 4 show a gradual convergence for both likelihood and impact with the largest distance equal to 0.2 in the first 100 iterations, where the variability of the data is also higher (i.e., the box size of the boxplots). The KS distance distribution becomes less variable and with lower median values after about 100 iterations (9% of the execution) and until 1500 iterations (68% of the execution). It is the case of mature partial results, that approximate the attack path features distribution from 15% to 5% (i.e., KS distances from 0.15 to 0.05). After 1500 iterations and until the end, we can recognize mature partial results with a KS distance very close to 0. In summary, attack path features can be queried after a few iterations with a 20% approximation, gradually decreasing reaching a 5% approximation after 1500 iterations, with a trend similar to vulnerability features (see Fig. 3).
@@
In the final validation step, we furnish evidence demonstrating that the attack path feature stability (Equation 3) accurately reflects the trend in statistical significance so that it can be used as a real-time indicator of the convergence to the GT. To this aim, Fig. 5 reports the trend of the stability for the performed experiments, where the trend of the median values is represented with full-color hue, while the alpha blended areas identify the variations between the upper and lower quartile values to report the statistical variability.
Figure 5: Attack path stability trend of the attack path features during the different iterations of the approach.
@@
The stability trend is definitely coherent with the KS distance trend of Fig. 4: during the first 100 iterations (9% of the execution), the stability indicates early partial results as well as the KS distance with the GT, with stability values lower than 0.85. Between iterations 100 and 1000 (46% of the execution), the stability reports an intermittent trend between 0.85 and 0.95, indicating the mature partial results, where the AG is assisting its statistical significance. The same trend corresponds to the KS distance from the GT. Finally, we can identify the definitive partial results in the last 54% of iterations, where the stability is very close to 1 and, correspondingly, the KS distance with the GT is very close to 0. In conclusion, the analysis of the stability indicates that it is a good predictor of the statistical significance of the partial AGs, with the great advantage of being a metric that can be computed in real-time for each iteration of the process.
@@
The StatAG validation showed the ability of the approach to generate the attack paths progressively, with three thresholds of significance (early, mature, and definitive) that inform about the convergence of the partial AG to the GT. This progressive approach represents an innovation in the Attack Graph community, where traditional solutions typically focus on the complete generation of AGs. However, when it comes to analyzing the complete AG for attack path analysis, this approach still requires enumerating all paths. To expedite convergence towards the subset of the complete AG that addresses a particular attack path query, the next section introduces a steering mechanism designed to guide both the generation and analysis processes.
5 SteerAG: Steered Attack Graph Generation and Analysis
@@
StatAG allows us to consider an approximation of the complete AG before the actual completion of its generation at every iteration of the process. This only partially addresses the original problem of the poor scalability of AG: while the convergence rate is the same as the random walk sampling, and different quality of results can be used for different analyses (e.g., exploration, confirmation), this may still require a long time for an analysis query to get the exact result. While for certain analyses this would not be needed, being sufficient working with a suitable approximation, for others this characteristic would not be ideal. For this reason, we contribute a further improvement of the framework, central in its capability to merge analysis and generation, and let the second be guided by the first, consisting of accelerating the attack path analysis for dynamic and real-time environments through the design of a steering mechanism, reported in Fig. 6. We name this approach Steered Attack Graph Generation and Analysis (SteerAG).
Figure 6: Workflow for the steered attack graph generation and analysis (SteerAG).
According to the general framework (see Fig. 1), the steering process is activated with the presence of attack paths generated using StatAG and an attack path query for the analysis, issued by a human user or an automatic process.
@@
The first step of the steering approach is the path features training, which consists of two activities. First, we label the attack paths coming from the StatAG as “relevant” when their attack path features correspond to the query requirements and “not relevant” otherwise. As this operation is linear in the number of attack paths generated up to the issuing of an analysis query in the worst case, it does not represent a costly operation. When a sufficient number of attack paths are labeled, with this number depending on the ML model used (i.e., ten to fifty for decision trees, around 100 for neural networks) the second activity is the actual training of a binary ML-based classification model, where the attack paths are classified according to their assigned label but the classification is based on the vulnerability features of the paths and not on characteristics of the whole path. In this way, we aim to discover the latent relations existing between all the relevant attack paths and their vulnerability characteristics based on the ones available. Being capable of extracting this relation allows for steering the AG generator by the discovered intervals of vulnerability features.
@@
In our proposal, we use the decision tree classifier [50] as the ML model for its properties that fit well the steering approach. Indeed, decision trees (i) can be trained quickly and therefore are suitable for progressive data analysis [33], (ii) produce sufficiently powerful classification models from relatively small inputs, allowing its use early on during the progression [50], and (iii) their structure can be easily translated into steering rules which represent the features of the relevant data [51].
@@
Once the decision tree has been trained on the vulnerability features, the goal of the vulnerability features extraction phase is to retrieve the rules to steer the next StatAG generation from the structure of the trained decision tree. We name them steering rules and must consider the values of the vulnerability features that generated only the “relevant” paths class. To do so, we consider that the decision tree is built by using the vulnerability features as split points and a leaf of the tree identifies whether the vulnerability features from the root to the leaf generate “relevant” paths or not. This makes it possible to compute the inverse mapping from the “relevant” leaves to the root, keeping track of the vulnerability features values that they and only they have in common. We can build the set of steering rules considering that each path from the root to a “relevant” leaf node represents a conjunction of decision rules that must be satisfied in order for the decision tree to classify it as relevant. Consequently, the logical disjunction of all such decision rules generates the set of steering rules. Depending on the depth of the decision tree, we could have many steering rules, where their union creates latent relations between relevant attack paths and vulnerability features.
@@
In the next step of the approach, we use the steering rules to consider only the vulnerabilities that match them before sampling the graph again with random walks as the StatAG generator mandates. In this way, the newly generated attack paths have a high probability of matching the “relevant” class for the attack path query given in input, thus accelerating the progression of the AG generation towards attack paths answering the query.
@@
Let us note that while the system is running, it continues to use the latest retrieved steering rules until there is a slowdown in the precision of the generation process, indicating that the relevant attack paths have been exhausted. To check that this effectively means having reached the exact answer to the query, another training is launched to adjust the steering with new rules that take into account the cumulative collection of attack paths, effectively repeating the described process until it converges to the exact result (i.e., a new training does not return any additional attack paths). At each iteration of SteerAG, preliminary attack path analysis can be performed to answer the query, until all paths are generated. The process is even faster in converging, guaranteeing an analysis results identical in quality to the one conducted on the fully generated AG.
In the rest of this section, we validate SteerAG, highlighting its performances by comparing it to the case in which it is not used.
@@ 5.1 SteerAG Validation
For the validation of SteerAG, we use the setting configurations used for StatAG and described in Section 4.1.1, with the addition of another experimental parameter that is a set of 100 attack path queries obtained with different combinations of attack path features in the query and varying their ranges from 0 to 1 with steps of 0.1. This resulted in a total number of 15,000 experiments.
@@
Figure 7: Comparison of (a) recall and (b) precision trends for StatAG (without steering) and SteerAG (with steering). The median values trend for each curve is reported with full-color hue, while the alpha blended areas identify the variations between the upper and lower quartile values.
Fig. 7 reports the recall and precision of the performed experiments. The median values trend for each curve is reported with full-color hue, while the alpha blended areas identify the variations between the upper and lower quartile values. In this way, the statistical variability is reported for all curves.
@@
From the recall trend (Fig. 7a) we can observe the very good performance of SteerAG that reaches values very close to 1 after just 250 iterations, with the activation of the steering mechanism starting at 100 iterations on average. In particular, we can recognize two thresholds during SteerAG: between iteration 100 and 150 we observe early partial results (recall lower than 0.5), useful for exploration; from iteration 200 until iteration 250 (11% of the execution) the results are mature, supporting early decision making; finally, from iteration 250 on we observe definitive partial results, with a median recall of more than 0.95 indicating that 95% of the relevant paths are retrieved.
@@
It is interesting to note that without the presence of the steering approach, the convergence to the GT (more precisely, to the portion of the GT relevant to the query) is very slow: at the definitive partial results threshold (iteration 250), random sampling has achieved a recall value of 0.2. Comparing the two trends, we can conclude that the steering approach accelerates the convergence to the GT saving around 90% iterations, providing a great advantage for the timely analysis of the attack paths. These performances are achieved consistently in all the experiments tested.
While the recall describes the ability of SteerAG to rapidly converge to the GT, the precision is used as a real-time indicator of the quality of each partial result. Fig. 7b shows the median precision trend during the execution of the framework.
@@
We can observe that the precision values are consistently high after the activation of the steering mechanism at iteration 100, presenting median values that reach peaks of 0.85. Similarly to the recall trend, precision presents a rapid growth in the early iterations, in particular during the generation of early partial results; this shows that the analyst can expect a fast transition from early to mature partial results, due to the high precision values. At iteration 500 (23% of the execution) it slows down. After 700 iterations (at 55% of the execution), there is another peak of precision up to 0.2, and this can also be observed in the recall trend in Fig. 7a. It is due to the retraining of the decision tree to adjust the steering rules after the detection of the precision breakdown. The fact that the precision has lower median values in the second steering activation than in the first one can be explained by a lower residual probability of finding the few relevant attack paths left (less than 2% at that stage).
In contrast, StatAG without the steering mechanism has a constant trend of low precision values because, as expected, the probability of uniformly randomly picking relevant attack paths to the issued query is much lower.
Overall, the validation of recall and precision demonstrate how the SteerAG accelerates the AG generation and analysis. While it gives the statistical variability of the trends, it does not allow us to investigate deeply how the characteristics of the attack path queries affect the approach performance. In particular, we are interested in evaluating the effect that the stringency of a query has on the performances, to avoid cases in which a stringent query, set to retrieve a low number of attack paths very specific, is indefinitely running due to slow convergence. To validate this, we define three ranges of queries: queries with low range when the values of the attack path features are in a range of at most 0.2, medium range when it is between 0.3 and 0.5, and high range when it is over 0.5. The rationale is that low queries require very specific characteristics of vulnerabilities with narrow intervals, and so are more specific. Large queries retrieve a large number of attack paths, defining not stringent vulnerability values in large intervals. Medium queries are in the middle. We report different analyses in Fig. 8.
(a) Recall trend.
(b) Convergence rate.
(c) Missing iterations from the one with recall 1.
Figure 8: Analysis of the steering mechanism in relation to the range of the attack path query.
Fig. 8a shows the recall trend of the different experiments for each query range. It highlights three clearly different trends that for the low, medium, and high ranges are respectively the fastest, medium, and slowest convergence to the GT. To quantify the speed of the convergence to the GT, in Fig. 8b we plot the recall values at the point of the maximum convergence of the recall curves. In practice, it corresponds to the peak of the recall curve before its stability. If the maximum convergence rate is low ($<0.2$) for high recall values ($>0.8$), then it means that the approach converges slowly to the GT. It is the case of 100% of the experiments with high query ranges, 54% of the experiments with medium ranges, and 25% of the ones with low query ranges. In contrast, for 46% of the experiments with medium ranges and 75% of the ones with low ranges, the convergence rate is high ($>0.3$) when the recall is higher than 0.8, indicating a very quick convergence to the GT. In particular, 25% of the experiments with low ranges reach the maximum convergence rate (between 0.2 and 0.3) when the recall is around 0.8, indicating a slower convergence with respect to the other low-range experiments.
Finally, in Fig. 8c we study the number of iterations remaining from the point of maximum convergence until all relevant paths are retrieved (i.e., when the recall is 1). We observe two main portions: the left one indicates the experiments that are closer to the recall value 1 and it is composed of all the experiments with low range and 70% of the ones with medium range. The right one indicates the cases that require more iterations to reach recall 1 and includes all the experiments with the high range and 30% of the ones with the medium range.
The results for low and medium query ranges confirm the capability of SteerAG to retrieve fast attack paths even for stringent queries correctly. The results for the high-range queries seem counter-intuitive to be not as good as the others. In fact, this is not true, as an investigation of the trend shows that it depends only on the very high quantity of attack paths to retrieve, related to the capacity to process attack paths at each iteration (i.e., the sampling size for each iteration) than to a difficulty of the approach in retrieving correct attack paths. Looking at the precision values, they are consistently high throughout the whole process, and it is worth noting that the maximum convergence rate is always reached with high recall values ($>0.8$), even for high-range queries.
Overall, we do not observe an effect of the stringency of the query on the performance, with the only parameter to set as the sampling size in order to speed up the high-range queries.

### v2

@@ 4.1 StatAG Validation
To validate StatAG we present the experimental setting, then we analyze the convergence of vulnerability and attack path features, and stability distributions.
@@
Experimental Setting. We validate StatAG with an experimental setup designed for approaching the different factors that affect the AG scalability [49]. It consists of synthetic networks and vulnerability inventories, in which we vary the number of hosts, vulnerabilities, network topology, and composition of vulnerabilities per host. Looking at the former, we considered configurations with up to 20 hosts and 50 vulnerabilities per host, which is the biggest configuration possible to compute all the paths (complete AG, used as ground truth) in a reasonable amount of time (in the order of hundreds of hours) [41]. We then tested $3$ types of network topologies, including mesh, random, and power law. We considered $5$ levels of heterogeneity of vulnerabilities, varying from 0% (i.e., all vulnerabilities among the hosts are the same) to 100% (i.e., all vulnerabilities among the hosts are different) to increase the variability of vulnerability features and $4$ vulnerability distributions (Uniform, Pareto, Binomial, and Poisson) to simulate realistic scenarios. Thus, for each network, we evaluated 60 different configurations, which we varied in 100 statistical variations for a total of 6000 experiments. We use the scikit-learn library [35] for the implementation and we run the experiments on a Linux server with Intel(R) Xeon(R) Gold 6248 CPU 2.50GHz and 256 GB memory.
@@
Results. We validate StatAG by studying the convergence of the partial AGs to the complete AG. For the sake of simplicity, in the rest of this section, we refer to the complete AG as ground truth (GT) to indicate that it corresponds to all the paths generated at the end of the classic generation process.
Figure 3: KS distance of partial AGs from the GT (vulnerability features).
@@
The first validation step concerns the convergence of the vulnerability feature distributions (from CVSS) of the partial AGs to the GT. Fig. 3 reports the median KS distances during the entire approach execution. The convergence speed of the vulnerability feature distributions depends on the distance threshold $T$ (see Definition 6). For example, considering a good approximation as $T=0.1$ (corresponding to just 10% approximation of the GT), 9 out of 10 features are statistically significant after just 100 iterations, except for the availability feature that has a KS distance of 0.2 after 100 iterations. In particular, we can notice three different trends in all the plots in Fig. 3. The first one coincides with early partial results, where the KS distance from the GT is too high to perform analysis with a reasonable approximation but suitable for initial exploration. This trend lasts about the first 100 iterations. After 100 iterations, corresponding to 9% of the complete execution, the trend is flat until the 1700-th iteration, corresponding to 77% of the complete execution.
Figure 4: KS distance of partial AGs from GT (attack path features).
This is the interval where mature partial results appear, that provide a 10% approximation of the GT for most vulnerability features, with the worst performance reached by the availability with a 20% approximation. Finally, in the last 23% of the execution, the KS distance becomes very close to 0, corresponding to the definitive partial results. Consequently, the vulnerability variability is captured in just 9% of the total iterations (the moment in which mature partial results appear), identifying the iteration from which an analyst can start reasoning on the AG. The convergence trend of the vulnerability features to the GT shows the capability of the progressive approach to represent 90% of the vulnerability inventory after a few iterations. It is a good result for the analysis of the vulnerability inventory, but we need further investigation that concerns the attack path analysis.
@@
To this aim, the next validation step concerns the convergence trend of the attack path features of the partial AGs to the GT ones. Fig. 4 reports the distribution trend of the KS distances between the partial AGs and the GT for the attack path features, i.e., likelihood and impact. The trends show a gradual convergence for both likelihood and impact, with the largest distance equal to 0.2 in the first 100 iterations, where the variability of the data is also higher (i.e., the box size of the boxplots). The KS distance distribution becomes less variable and with lower median values after about 100 iterations (9% of the execution) and until 1500 iterations (68% of the execution). It is the case of mature partial results, that approximate the attack path features distribution from 15% to 5% accuracy (i.e., KS distances from 0.15 to 0.05). After 1500 iterations and until the end, we can recognize definitive partial results with a KS distance very close to 0. In summary, attack path features can be queried after a few iterations with a 20% approximation, gradually decreasing, reaching a 5% approximation after 1500 iterations, comparable to vulnerability features (Fig. 3).
@@
In the final validation step, we furnish evidence demonstrating that the attack path feature stability (Equation 2) accurately reflects the trend in statistical significance so that it can be used as a real-time indicator of the convergence to the GT. To this aim, Fig. 5 reports the stability trend for the performed experiments, where the trend of the median values is represented with full-color hue, while the alpha blended areas identify upper and lower quartile values to report the statistical variability.
Figure 5: Stability for the path features.
@@
The stability trend is coherent with the KS distance trend of Fig. 4: during the first 100 iterations (9% of the execution), the stability indicates early partial results as well as the KS distance with the GT, with stability values lower than 0.85. Between iterations 100 and 1000 (46% of the execution), the stability reports an intermittent trend between 0.85 and 0.95, indicating the mature partial results, where the AG is assisting its statistical significance. The same trend corresponds to the KS distance from the GT. Finally, we can identify the definitive partial results in the last 54% of iterations, where the stability is very close to 1 and, correspondingly, the KS distance with the GT is very close to 0.
@@
In conclusion, this validation experimentally proves the correlation between stability and statistical significance in different configurations. While statistical significance can be calculated only a posteriori, stability has the great advantage of being computed in real-time, thus informing the analyst at each iteration. The StatAG validation showed the ability of the approach to generate the attack paths progressively, with three thresholds of significance (early, mature, and definitive) that inform about the convergence of the partial AG to the GT. This progressive approach represents an innovation in the Attack Graph community, where traditional solutions typically focus on the complete generation of AGs. It allows continuously fed data from the AG at every iteration and is more efficient in computing a statistically representative version of the AG. However, when it comes to analyzing the complete AG for attack path analysis, this approach still requires enumerating all paths. To expedite convergence towards the subset of the complete AG that addresses a particular attack path query, the next section introduces a steering mechanism designed to guide both the generation and analysis processes.
5 SteerAG: Steered Generation and Analysis
Figure 6: SteerAG workflow.
@@ 5 SteerAG: Steered Generation and Analysis
StatAG provides an approximation of the entire AG during each iteration, only partially solving the scalability issue, as exact results may still take a long time to obtain for some queries due to low convergence speed despite offering varying result qualities for different analyses. For this reason, we contribute a further improvement of the framework, central in its capability to merge analysis and generation, and let the second be guided by the first, consisting of accelerating the attack path analysis for dynamic and real-time environments through the design of a steering mechanism, reported in Fig. 6. We name this approach Steered Attack Graph Generation and Analysis (SteerAG). According to the framework (see Fig. 1), the steering process is activated with the presence of partial attack paths (PAP in Fig. 6) generated using StatAG and an attack path query for the analysis, issued by a human user or automatic process.
@@
The first step of the steering approach is the path features training, which consists of two activities. First, we label the attack paths coming from the StatAG as “relevant” when their attack path features correspond to the query requirements and “not relevant” otherwise. As this operation is linear in the number of attack paths generated up to the issuing of an analysis query in the worst case, it does not represent a costly operation. When a sufficient number of attack paths are labeled, with this number depending on the ML model used (i.e., ten to fifty for decision trees, around 100 for neural networks), the second activity is the actual training of a binary ML-based classification model, where the attack paths are classified according to their assigned label. The classification is based on the vulnerability features of the paths and not on the characteristics of the whole path. In this way, we aim to discover the latent relations between all the relevant attack paths and their vulnerability characteristics based on the ones available. Extracting these relations allows for steering the AG generator by the discovered intervals of vulnerability features.
@@
In our proposal, we use the decision tree classifier [21] as the ML model for its properties that fit well with the steering approach as they: (i) can be trained quickly and therefore are suitable for progressive data analysis [13], (ii) produce sufficiently powerful classification models from relatively small inputs, allowing its use early on during the progression [21], and (iii) can be easily translated into steering rules, which represent the features of the relevant data [4].
@@
Once the decision tree has been trained on the vulnerability features, the goal of the vulnerability features extraction phase is to retrieve the rules to steer the next StatAG generation from the structure of the trained decision tree. We name them steering rules and must consider the values of the vulnerability features that generated only the “relevant” paths class. To do so, we consider that the decision tree is built by using the vulnerability features as split points, and a leaf of the tree identifies whether or not the vulnerability features from the root to the leaf generate “relevant” paths. This makes it possible to compute the inverse mapping from the “relevant” leaves to the root, keeping track of the vulnerability features values that only they have in common. We can build the set of steering rules considering that each path from the root to a “relevant” leaf node represents a conjunction of decision rules that must be satisfied for the decision tree to classify it as relevant. Consequently, the logical disjunction of all such decision rules generates the set of steering rules.
@@
In the next step of the approach, we use the steering rules to consider only the vulnerabilities that match them before sampling the graph again with random walks, as the StatAG generator mandates. In this way, the newly generated attack paths have a high probability of matching the “relevant” class for the attack path query given in input, thus accelerating the progression of the AG generation towards attack paths answering the query.
@@
Let us note that while the system is running, it continues to use the latest retrieved steering rules until there is a slowdown in the precision of the generation process, indicating that the relevant attack paths have been exhausted. To check that this effectively means having reached the exact answer to the query, another training is launched to adjust the steering with new rules that take into account the cumulative collection of attack paths, effectively repeating the described process until it converges to the exact result (i.e., a new training does not return any additional attack paths). At each iteration of SteerAG, preliminary attack path analysis can be performed to answer the query, until all paths are generated. The process is even faster in convergence, guaranteeing analysis results that are identical in quality to the ones conducted on the fully generated AG.
@@ 5.1 SteerAG Validation
For the validation of SteerAG, we use the setting configurations used for StatAG and described in Section 4.1, with the addition of another experimental parameter that is a set of 100 attack path queries obtained with different combinations of attack path features in the query and varying their ranges from 0 to 1 with steps of 0.1. This resulted in a total number of 60,000 experiments.
@@
Figure 7: Comparison of (a) recall and (b) precision for StatAG and SteerAG. Median values are with full-color hue, and alpha blended areas indicate quartiles.
@@
Fig. 7 reports the recall and precision of the performed experiments. The median values trend for each curve is reported with full-color hue, while the alpha blended areas identify the variations between the upper and lower quartile values. From the recall trend (Fig. 7a) we can observe the very good performance of SteerAG that reaches values very close to 1 after just 250 iterations, with the activation of the steering mechanism starting at iteration 100 on average. In particular, we can recognize two thresholds during SteerAG: between iteration 100 and 150 we observe early partial results (recall lower than 0.5 but fastly increasing), useful for exploration; from iteration 200 until iteration 250 (11% of the execution) the results are mature (recall higher than 0.9), supporting early decision making; finally, from iteration 250 on we observe definitive partial results (recall higher than 0.98).
@@
It is interesting to note that without the steering approach, the convergence to the GT is very slow: at the definitive partial results threshold (iteration 250), random sampling has achieved a recall value of just 0.2. Comparing the two trends, we can conclude that the steering approach accelerates the convergence to the GT, saving around 90% of the iterations consistently in all the experiments tested, providing a more timely analysis of attack paths.
@@
While the recall describes the ability of SteerAG to rapidly converge to the GT, the precision informs the analysts about the percentage of attack paths answering the query at each iteration: thus, high values correspond to many paths answering the query. Fig. 7b shows the median precision trend during the framework execution. We can observe that the precision values are consistently high after the activation of the steering mechanism at iteration 100, presenting median values that reach peaks of 0.85. Similarly to the recall trend, precision presents rapid growth in the early iterations, particularly during the generation of early partial results; this shows that the analyst can expect a fast transition from early to mature partial results. At iteration 500 (23% of the execution) it slows down. After 700 iterations (at 55% of the execution), there is another peak of precision up to 0.2 (visible also in the recall trend in Fig. 7a. It is due to the retraining of the decision tree to adjust the steering rules after the detection of the precision breakdown. Indeed, a precision breakdown happens only when most of the paths are already retrieved, aligning to low values typical of random sampling (end of steering phase). The fact that the precision has lower median values in the second steering activation than in the first one can be explained by a lower residual probability of finding the few relevant attack paths left (less than 2% at that stage). In contrast, StatAG without the steering mechanism has a constant trend of low precision values because, as expected, the probability of uniformly randomly picking relevant attack paths to the issued query is much lower. Finally, the performance is not significantly affected by the stringency of the query (i.e., different query ranges) as we experimentally proved in Appendix 0.A.